You are a shareholder in a C corporation. The corporation earns$ 1.85per share before taxes. Once it has paid taxes it will distribute the rest of its earnings to you as a dividend. Assume the corporate tax rate is 40% and the personal tax rate on​ (both dividend and​ non-dividend) income is 30%.

How much is left for you after all taxes are​ paid?

You are a shareholder in an S corporation. The corporation earns $ 1.63per share before taxes. As a pass through​ entity, you will receive $ 1.63 for each share that you own. Your marginal tax rate is20%.

How much per share is left for you after all taxes are​ paid?
Answer #1

Answer 1st part:

C Corporation earns $1.85 per share before taxes.

Corporate tax rate is 40%

Once C Corporation has paid taxes it will distribute the rest of its earnings to shareholder as a dividend.

Dividend received = $1.85 * (1 - 40%) =$1.11 per share.

Personal tax rate on​ (both dividend and​ non-dividend) income is 30%

Hence:

Amount left to shareholder after all taxes are​ paid = $1.11 * (1 - 30%) = $0.777 per share

Amount left after all taxes are​ paid = $0.777 per share

Answer 2nd part:

S corporation is a pass through entity and it does not pay taxes.

The corporation earns $ 1.63 per share before taxes.

As a pass through​ entity, shareholder will receive $ 1.63 for each share that he/she owns.

Shareholders' marginal tax rate is 20%.

Amount per share is left after all taxes are​ paid = $1.63 * (1 - 20%) = $1.304

Amount per share is left after all taxes are​ paid = $1.304